OpenStudy (anonymous):

Maria has three children.There is two years age difference between each child.The total ages of three children is 36 years.Tina is the youngest child.How old is Tina let T=the age of Tina

OpenStudy (anonymous):

t is tina t+2 is another one t+4 is the biggest child 3t+6=36 3t=30 t=5

OpenStudy (hoblos):

T=the age of Tina the ages of the other two children are (T+2) and (T+4) thus T + (T+2) +(T+4)=36 3T +6 = 36 3T = 30 T=10 years

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

@strenesmee 30/3 = 10 not 5 .. ! @eruiz u r welcome :)

OpenStudy (anonymous):

yeah! i made a silly mistake

OpenStudy (hoblos):

it always happen :P